OpenStudy (anonymous):

17x - 6 + 3x - 5 = x + 11 + 4x

OpenStudy (anonymous):

need help please

OpenStudy (anonymous):

20x - 11 = 5x + 11 20x + 1 = 5x + 11 9x = 16x

OpenStudy (cggurumanjunath):